# When LLMs Play the Telephone Game: Cumulative Changes and Attractors in Iterated Cultural Transmissions

### v1

@@ 1 Introduction
Large language models (LLMs) are playing an increasingly significant role in the production of media content across various domains [10]. They are being used for academic writing [12, 2, 16, 24, 37, 40, 14], journalism [55], story generation [70, 78, 79], as chatbots on social media [61] and in the workplace at large [11, 27]. As LLMs become more performant, controllable, and more widespread, their impact on the creation and dissemination of information content humans consume is expected to grow even further [10].
@@
Given the implications of LLM usage for the production of cultural information, numerous researchers have studied the properties of LLM-generated content. In these studies, LLMs showed several biases with respect to gender [1, 62, 72, 30, 25], race [57], values [5, 66], politics [48, 3, 30], authority, fallacy oversight, and beauty [15]. They were also found to generate at least as attractive [40] and compelling [68] texts than humans and to display similar cognitive biases [26].
@@
While single-turn behaviors of LLMs prompted with a human-generated prompt are under active scrutiny, little is known about the effect of iterated interactions. Indeed, LLMs often use existing cultural content to generate new ones, for example when writing scientific reviews [37, 2] or generating new stories based on existing examples [78]. As the share of AI-generated content increases, LLMs will be producing outputs using content that is already LLM-generated, making it crucial to study the consequence of this iterative process. Moreover, LLMs are increasingly being used in multi-agent settings [20, 51, 52, 77, 34, 71, 17] and are already interacting with one another as chatbots on social media111https://chirper.ai/.
@@
Nonetheless, little is known about how populations of LLMs might self-organize, as other complex systems do. Research in complex systems traditionally studies how global-level patterns emerge from local interactions [29, 43]. Here, we ask whether multi-turn behaviors conditioned on LLM-generated content cause the appearance of new kinds of biases, undetectable in single-turn behaviors but accumulating across iterated interactions.
To address this question, we take inspiration from the cultural evolution literature. The field of cultural evolution aims to provide causal explanations for the change of culture (defined as socially inherited information) over time. In particular, we draw insights from a research tradition called cultural attraction theory (CAT) [67, 47, 44]. CAT aims to determine how non-random transformations of cultural information during transmission events may lead to the evolution of progressively more stable forms, referred to as attractors. Although the precise conceptualization of attractors varies across authors, an encompassing definition may be "theoretical posits that capture the way in which certain ideational variants are more likely to be the outcome of transformations than others." [13]. For example, CAT has shown how the cognitive appeal of bloodletting — the historically commonplace but often damaging practice of letting blood out to cure a patient — explains why it is found in many unrelated cultures worldwide [45]. Experiments from the CAT literature showed that successive transmissions of a story about a mundane event can lead bloodletting to acquire a causal role that it did not have in the original story. Other experimental work has also revealed how modifying information over transmission chains makes it converge toward inductive biases [36] or how ecological factors may influence the position of attractors [46].
@@
Most experiments in CAT employ a transmission chain design [42], first introduced by [7]. In this setup, chains of participants receive, produce and transmit social information from and to each other in a sequential manner (as in the popular telephone game). This powerful and highly controlled design allows to evaluate the high-level patterns that emerge from the accumulation of directional changes during single-turn transmission events. Here, we adapt this design to study how culture evolves along chains of LLMs, rather than human participants (Figure 1).
@@
We conducted several transmission chain experiments with LLMs, where the first LLM-based agent in the chain receives a human-generated text, elaborates on it, and then passes it to the next agent in the chain. This transmission step is repeated with different instances of the LLM agent until the end of the chain is reached. By introducing several novel evaluation methods, we estimate the extent to which successive transmission events affect the evolution of multiple text properties, namely its toxicity, positivity, difficulty, and length. By comparing the properties of the initial (human-generated) and final texts (after several transmissions between LLMs), we illustrate and study the existence of potential attractors in LLM cultural evolution. In particular, we measure the effect of consecutive interactions compared to the effect of the first one. To study how text properties and attractors are affected by the specific context in which culture evolves, we conduct our analyses on five different models (ChatGPT-3.5-turbo-0125, Llama3-8B-Instruct, Mistral-7B-Instruct-v0.2, Llama3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1), three different tasks (i.e., instructions to either “rephrase”, “take inspiration from”, or “continue” the initial text) and 20 different initial texts. Although our focus is on a better understanding of the cultural dynamics of LLMs, the metrics and evaluation methods introduced here may also be of interest to researchers studying human cultural evolution.
@@
The code for reproducing the simulations, analyses and figures is available on our GitHub222https://github.com/jeremyperez2/TelephoneGameLLM.
@@ 2 Related work
LLM-generated content is known to exhibit a variety of stereotypical biases [8, 73]. In single-turn settings, LLMs perpetrate [49] or even amplify human biases based on gender, nationality, race, and religion [39]. For instance, the GPT model was shown to exhibit cultural values similar to those of WEIRD (Western, Educated, Industrial, Rich, Democratic) cultures [5]. LLMs trained through reinforcement learning with human feedback (RLHF) were found to overly express left-wing opinions on American politics — a tendency that, once formed, is difficult to avoid even after steering the model toward different demographic groups [63].
@@
Several studies have applied experimental designs used in cultural evolution to study knowledge and skill accumulation in groups of Reinforcement Learning agents [18, 64, 69, 56]. Closer to the current study, populations of LLMs have also been studied [10]. Iterative chains of generative models trained on the preceding model’s output have been shown to sometimes collapse toward the most likely outputs while the tails of the original distribution disappear [65, 54]. This idea of using LLM-generated content to fine-tune the next generation has also been applied to groups of LLMs with various communication structures [33]. Similar to our approach, iterative chains with frozen (i.e., not re-trained) LLMs have been shown to express human-like biases in terms of gender stereotypes, positivity, and social, threat, and biology-related information [1]. Strong, but non-human-like biases for producing factual information have also been observed [17], stressing the importance of understanding the evolution of content in LLMs and the ways it might deviate from human cultural evolution.
@@ 3 Methods
Our telephone game experiments aim to study the possible attractors and biases that may accumulate across multiple turns of interactions between LLMs. This is done with a transmission chain design tracking the evolution LLM outputs as a function of the number of interactions in the LLM chain. This section introduces our transmission chain design (Section 3.1), the set of metrics used to study the evolution of text properties, semantic similarity and the added effect of multi-turn interactions (Section 3.2), and our method to characterize the properties of attractors (Section 3.4).
@@ 3.1 LLM transmission chains
where text0 is the initial human-generated text and generates an output based on task and the previous agent’s text . We run this process for 50 generations. Examples of texts evolving through generations are provided in Appendix Section A.
@@
We borrow human-generated text from various databases to provide the initial input to each transmission chain. Since we were interested in how variation in the initial text would impact the properties of the ensuing chain, human-generated texts spanned various types of content: scientific abstracts 333https://huggingface.co/datasets/CCRss/arxiv_papers_cs, news articles 444https://huggingface.co/datasets/RealTimeData/bbc_latest, and social media posts 555https://huggingface.co/datasets/FredZhang7/toxi-text-3M/blob/e0e5b168b4a7e14e84f07271bfe1c6b42bc91ccd/multilingual-train-deduplicated.csv. As we are interested in the evolution of the toxicity, positivity, difficulty and length of generated texts, we sample the entire dataset to obtain a subset of 20 initial texts that covered the range of possible values for these properties. The exact method used to extract these texts is detailed in Appendix Section A.
@@
As for models, tasks remained consistent within each chain. The exact prompt used for each task is reported in Appendix Section A.
@@
To assess whether and how cultural evolution dynamics are affected by the model specifications, we run identical experiments using five different models, all commonly used, from three different companies and with varying sizes: GPT-3.5-turbo-0125 (referred to as GPT3.5) , Llama3-8B-Instruct (referred to as Llama3-8B), Mistral-7B-Instruct-v0.2 (referred to as Mistral-7B), Llama3-70B-Instruct (referred to as Llama3-70B), Mixtral-8x7B-Instruct-v0.1 (referred to as Mixtral-8x7B). For inference, we used the OpenAI API (The MIT License) 666https://openai.com/index/openai-api/ to run GPT3.5 and the HuggingFace’s Transformer library [76] for other models (Apache Licence, v2.0). In our setup, transmission chains are always homogeneous with respect to the model, i.e. each chain is composed of a population of agents sharing the same underlying model.
@@ 3.2 Metrics
Toxicity. Companies typically fine-tune LLMs to avoid the generation of toxic (i.e., dangerous or harmful) content. However, to our knowledge, this fine-tuning step focuses on single-turn dynamics, and the evolution of content with respect to its toxicity is currently understudied. We measure the toxicity of texts by quantifying the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in [31].
Positivity. Even when trained to avoid toxic content, LLMs have been shown to express similar positivity biases to humans, often favoring negative over positive information in preserving and generating new information [1]. To study whether positivity biases over transmission chains are affected by tasks and models, we measure the positivity of produced contents using the SentimentIntensityAnalyzer tool from NLTK [32]. It uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
Difficulty. While LLMs are argued to benefit society by democratizing knowledge [74], such positive outcomes are conditioned on the LLMs generating output that is accessible and inclusive to all kinds of audiences. However, whether text difficulty is preserved, increased, or reduced over transmission chains is currently unknown. We estimate text difficulty using the Gunning-Fog index [9], which depends on the average sentence length and the percentage of difficult words. A standard interpretation of this index is that it estimates the years of formal education required to fully understand the text.
@@
We provide additional details about metrics in Appendix Section A.
Semantic similarity
As we are also interested in measuring the extent to which different chains converge (or diverge) towards producing semantically similar (or different) outputs, we measure the similarity between all pairs of text produced. We do so by computing the cosine similarity between text embeddings, obtaining a value of similarity in the range . An average similarity score close to 1 indicates all texts were similar to each other, whereas scores close to 0 and below indicates pairs of texts were highly dissimilar. Computing similarity scores for text outputs requires text embeddings, which we obtained with an embedding model from Sentence-Transformer [59].
@@ 3.3 Effect of multi-turn transmissions
One of our questions is how content evolves over multi-turn transmissions compared to single-turn settings. To address this point, we compare the distribution of a given property in the generated texts at the first generation to the distributions at subsequent generations. Thus for each model and task, we look at the properties of each of the 100 generated texts (20 transmissions chains * 5 seeds) at each generation, which gives us a sample of 100 property values for each value. Using a Kolmogorov–Smirnov test [41], we then test whether the sample obtain at each generation comes from the same distribution as the sample obtained after the first generation. If we can confidently reject the hypothesis that the sample of property values at the end of the transmission chain comes from the same distribution as the sample obtained after the first generation, this would confirm that looking at outputs after a single-turn transmission is not enough for predicting output properties in a multi-turn setting.
@@ 3.4 Attractor strength and position
Human cultural evolution shows that cultural traits sometimes evolve towards attractor states, i.e., content that invites convergence even with different starting points [36, 45, 46, 13]. Therefore, we were interested in whether transmission chains with LLMs would show similar attractor dynamics, and whether these depend on the model and task used in the chain. The concept of cultural attractor is not consistently formalized in the human cultural evolution literature [13]. Here, we defined attractors as the theoretical equilibrium point to which the iterated generation process (defined in Eq. 1) may eventually converge. We mathematically define attractors in terms of two properties of interest: its position (i.e., the location in output generation space the process converges toward) and its strength (i.e., the intensity to which generated outputs are pulled toward it). The strength takes values in , which allows for a continuous notion of an attractor: rather than being a binary concept that either exists or does not, attractors here lie on a spectrum, covering systems without attraction effects (strength=0) to ideal attractors (strength=1). To compute position and strength, we use the simulated data to fit a linear regression predicting the value of a property at the end of the chain as a function of its value in the initial text (Figure 2).
@@
where I is the estimated intercept and s the estimated slope.
This enables us to estimate the final output of a new chain starting from the final output of the previous chain as:
@@
where .
If , then the sequence converges, its limit is and its convergence rate is .
We can therefore use the estimated relationship Eq. 2 to determine if an attractor exists () and, if so, estimate its position and strength .
To validate that these theoretical fixed points correctly capture attraction dynamics, we estimated their positions using only data from the first 10 generations of each chain, and compared the predictions with the actual output after 50 generations. Visual inspection of the results (Appendix Section B) confirmed that our method is suited for estimating the strength and position of attractors.

### v2

@@ 1 Introduction
Large language models (LLMs) are playing an increasingly significant role in the production of media content across various domains machine_culture. They are being used for academic writing buruk_academic_2023; agarwal_litllm_2024; cheng_have_2024; dergaa_human_2023; khraisha_can_2024; marlow_ghost_2021; cabanac_tortured_2021, journalism petridis_anglekindling_2023, story generation valentini_automatic_2023; xie_next_2023; zhao_more_2023, as chatbots on social media sadikoglu_evolution_2023 and in the workplace at large brynjolfsson_generative_2023; eloundou_gpts_2023. As LLMs become more performant, controllable, and more widespread, their impact on the creation and dissemination of information content humans consume is expected to grow even further machine_culture.
@@
Given the implications of LLM usage for the production of cultural information, numerous researchers have studied the properties of LLM-generated content. In these studies, LLMs showed several biases with respect to gender acerbi2023large; salewski_-context_2024; wan_kelly_2023; haller_opiniongpt_2023; dong_disclosure_2024, race raj_breaking_2024, values which_humans; sivaprasad_exploring_2024, politics motoki_more_2023; agiza_analyzing_2024; haller_opiniongpt_2023, authority, fallacy oversight, and beauty chen_humans_2024. They were also found to generate texts at least as attractive marlow_ghost_2021 and compelling spitale_ai_2023 as human-generated ones and to display similar cognitive biases echterhoff_cognitive_2024.
@@
While single-turn behaviors of LLMs prompted with a human-generated prompt are under active scrutiny, little is known about the effect of iterated text interactions. Indeed, LLMs often use existing cultural content to generate new material, for example when writing scientific reviews khraisha_can_2024; agarwal_litllm_2024 or generating new stories based on existing examples xie_next_2023. As the share of AI-generated content increases, LLMs will be producing outputs using content that is already LLM-generated, making it crucial to study the consequences of this iterative process. Moreover, LLMs are increasingly being used in multi-agent settings de_zarza_emergent_2023; park_generative_2023; park_social_2022; xiao_simulating_2023; hua_war_2023; vezhnevets_generative_2023; chuang2023simulating and are already interacting with one another as chatbots on social media111https://chirper.ai/.
@@
Nonetheless, little is known about how populations of LLMs might self-organize, as other complex systems do. Research in complex systems traditionally studies how global-level patterns emerge from local interactions gleick_chaos_1988; mitchell_complexity_2009. Here, we ask whether multi-turn behaviors conditioned on LLM-generated content cause the appearance of new kinds of biases, undetectable in single-turn behaviors but accumulating across iterated interactions.
To address this question, we take inspiration from the cultural evolution literature. The field of cultural evolution aims to provide causal explanations for the change of culture over time. In using the term culture, we borrow a definition traditionally used in the cultural evolution literature, where it refers to any type of socially transmitted information mesoudi_cultural_2016. This definition includes various types of cultural information (e.g. beliefs, skills, norms) and is independent of the medium of transmission (e.g. bodily imitation, oral or written transmission). Importantly, this way of defining culture is quite different from the view that is adopted by research works studying cultural values, or cultural alignment in LLMs, where culture is seen as an attribute of a specific human group (e.g. “Western culture”). This definition allows to use the same conceptual and methodological framework which was applied to study human derex_causal_2019 and non-human whiten_culture_2011 cultural evolution to study LLMs, which is why we adopt that definition for this work.
In particular, we draw insights from a research tradition called cultural attraction theory (CAT) sperber_anthropology_1985; morin_how_2016; miton_cultural_2024. CAT aims to determine how non-random transformations of cultural information during transmission events may lead to the evolution of progressively more stable forms, referred to as attractors. Although the precise conceptualization of attractors varies across authors, an encompassing definition may be “theoretical posits that capture the way in which certain ideational variants are more likely to be the outcome of transformations than others.” buskell_what_2017. Importantly, the existence of attractors would imply that for a broad range of starting conditions, different variants of a given cultural trait will converge towards the same state, that is, to a cultural trait that possesses specific properties.
@@
Most experiments in CAT employ a transmission chain design mesoudi_experimental_2021, first introduced by bartlett_remembering_1932. In this setup, chains of participants receive, produce and transmit social information from and to each other in a sequential manner (as in the popular telephone game). This powerful and highly controlled design allows to evaluate the high-level patterns that emerge from the accumulation of directional changes during single-turn transmission events. Here, we adapt this design to study how culture evolves along chains of LLMs, rather than human participants (Figure 1).
We conducted several transmission chain experiments with LLMs, where the first LLM-based agent in the chain receives a human-generated text, elaborates on it, and then passes it to the next agent in the chain. This transmission step is repeated with different instances of the LLM agent until the end of the chain is reached. We simulated repeated transmission to observe the direction in which text properties evolve, starting from a broad range of initial values. Finding that certain cultural variants (for example, texts with low toxicity) are more likely to be the outcome of this cultural evolutionary process would reveal the existence of an attractor.
@@
By introducing several novel evaluation methods, we estimate the extent to which successive transmission events affect the evolution of multiple text properties, namely its toxicity, positivity, difficulty, and length. By comparing the properties of the initial (human-generated) and final texts (after several transmissions between LLMs), we illustrate and study the existence of potential attractors in LLM cultural evolution. In particular, we measure the effect of consecutive interactions compared to the effect of the first one. To study how text properties and attractors are affected by the specific context in which culture evolves, we conduct our analyses on five different models (ChatGPT-3.5-turbo-0125, Llama3-8B-Instruct, Mistral-7B-Instruct-v0.2, Llama3-70B-Instruct, and Mixtral-8x7B-Instruct-v0.1), three different tasks (i.e., instructions to either “rephrase”, “take inspiration from”, or “continue” the initial text) and 20 different initial texts. Although our focus is on a better understanding of the cultural dynamics of LLMs, the metrics and evaluation methods introduced here may also be of interest to researchers studying human cultural evolution.
@@
The code for reproducing the simulations, analyses and figures is available on our GitHub222https://anonymous.4open.science/r/TelephoneGameLLM-D9E1.
@@
We show how the LLM properties (model and size) and the task influence the position and strength of attractors, as well as the consequence of fine-tuning and temperature tuning (Section 4.3).
@@ 2 Related work
LLM-generated content is known to exhibit a variety of stereotypical biases dangers_parrots; weidinger2021ethical. In single-turn settings, LLMs perpetrate nadeem2020stereoset or even amplify human biases based on gender, nationality, race, and religion kotek2023gender. For instance, the GPT model was shown to exhibit cultural values similar to those of WEIRD (Western, Educated, Industrial, Rich, Democratic) cultures which_humans. LLMs trained through reinforcement learning with human feedback (RLHF) were found to overly express left-wing opinions on American politics — a tendency that, once formed, is difficult to avoid even after steering the model toward different demographic groups santurkar2023whose.
@@
Several studies have applied experimental designs used in cultural evolution to study knowledge and skill accumulation in groups of Reinforcement Learning agents cook_artificial_2024; schmitt_kickstarting_2018; open_ended_learning_team_open-ended_2021; prystawski_cultural_2023. Closer to the current study, populations of LLMs have also been studied machine_culture. Iterative chains of generative models trained on the preceding model’s output have been shown to sometimes collapse toward the most likely outputs while the tails of the original distribution disappear shumailov2023curse; peterson2024ai. The idea of using LLM-generated content to fine-tune the next generation has also been applied to groups of LLMs with various communication structures helm_tracking_2024. Similar to our approach, iterative chains with frozen (i.e., not re-trained) LLMs have been shown to express human-like biases in terms of gender stereotypes, positivity, and social, threat, and biology-related information acerbi2023large. Strong, but non-human-like biases for producing factual information have also been observed chuang2023simulating, stressing the importance of understanding the evolution of content in LLMs and the ways it might deviate from human cultural evolution.
Iterated learning
Parallel works (ren_language_2024; zhu_eliciting_2024) have studied similar questions using different methodologies, namely the framework of Bayesian Iterated Learning (BIL) (kirby_iterated_2014; griffiths_language_2007). Under this framework, attraction is conceptualized as convergence towards priors, whereas we conceptualize it as the fixed point of a recurrent sequence. Therefore, BIL allows to make prediction about how an agents’ hypothesis about a given phenomenon will evolve through social interaction, while our method captures how a specific cultural trait will evolve. Moreover, our framework enables us to avoid assuming that LLMs possess “priors” in the same way as humans do, which might be problematic (see kovac_large_2023; shanahan_role-play_2023).
@@ 3 Methods
Our telephone game experiments aim to study the possible attractors and biases that may accumulate across multiple turns of interactions between LLMs. This is done with a transmission chain design tracking the evolution LLM outputs as a function of the number of interactions in the LLM chain. This section introduces our transmission chain design (Section 3.1), the set of metrics used to study the evolution of text properties (Section 3.2) the effects of multi-turn interactions 3.3, and our method to characterize the properties of attractors (Section 3.4).
@@ 3.1 LLM transmission chains
where text0 is the initial human-generated text and generates an output based on task and the previous agent’s text . We run this process for 50 generations. Examples of texts evolving through generations are provided in Appendix A.
@@
We borrow human-generated text from various databases to provide the initial input to each transmission chain. Since we were interested in how variation in the initial text would impact the properties of the ensuing chain, human-generated texts spanned various types of content: scientific abstracts 333https://huggingface.co/datasets/CCRss/arxiv_papers_cs, news articles 444https://huggingface.co/datasets/RealTimeData/bbc_latest, and social media posts 555https://huggingface.co/datasets/FredZhang7/toxi-text-3M/blob/e0e5b168b4a7e14e84f07271bfe1c6b42bc91ccd/multilingual-train-deduplicated.csv. As we are interested in the evolution of the toxicity, positivity, difficulty and length of generated texts, we sample the entire dataset to obtain a subset of 20 initial texts that covered the range of possible values for these properties. The exact method used to extract these texts is detailed in Appendix A.
@@
Tasks remained consistent within each chain. The exact prompt used for each task is reported in Appendix A.
@@
To assess whether and how cultural evolution dynamics are affected by the model specifications, we run identical experiments using five different models, all commonly used, from three different companies and with varying sizes: GPT-3.5-turbo-0125 (GPT3.5) , Llama3-8B-Instruct (Llama3-8B), Mistral-7B-Instruct-v0.2 (Mistral-7B), Llama3-70B-Instruct (Llama3-70B), Mixtral-8x7B-Instruct-v0.1 (Mixtral-8x7B). For inference, we used the OpenAI API (MIT License) 666https://openai.com/index/openai-api/ to run GPT3.5 and the HuggingFace’s Transformer library wolf2019huggingface for other models (Apache Licence, v2.0). In our setup, transmission chains are always homogeneous with respect to the model, i.e. each chain is composed of a population of agents sharing the same underlying model.
@@ 3.2 Metrics
Toxicity. Companies typically fine-tune LLMs to avoid the generation of toxic (i.e., dangerous or harmful) content. However, to our knowledge, this fine-tuning step focuses on single-turn dynamics, and the evolution of content with respect to its toxicity is currently understudied. We measure the toxicity of texts by quantifying the presence of rude, disrespectful, or unreasonable language, using a probability score that ranges from 0.0 (benign and non-toxic) to 1.0 (highly likely to be toxic), as estimated by the classifier introduced in Detoxify.
Positivity. Even when trained to avoid toxic content, LLMs have been shown to express similar positivity biases to humans, often favoring negative over positive information in preserving and generating new information acerbi2023large. To study whether positivity biases over transmission chains are affected by tasks and models, we measure the positivity of produced contents using the SentimentIntensityAnalyzer tool from NLTK hardeniya2016natural. It uses this information to calculate a sentiment score for the text, ranging from -1.0 (highly negative) to 1.0 (highly positive).
Difficulty. While LLMs are argued to benefit society by democratizing knowledge techstrongGenerativeNext, such positive outcomes are conditioned on the LLMs generating output that is accessible and inclusive to all kinds of audiences. However, whether text difficulty is preserved, increased, or reduced over transmission chains is currently unknown. We estimate text difficulty using the Gunning-Fog index bogert1985defense, which depends on the average sentence length and the percentage of difficult words. A standard interpretation of this index is that it estimates the years of formal education required to fully understand the text.
@@
We provide additional details about metrics in Appendix A.
@@ 3.3 Effect of multi-turn transmissions
One of our questions is how content evolves over multi-turn transmissions compared to single-turn settings. To address this point, we compare the distribution of a given property in the generated texts at the first generation to the distributions at subsequent generations. Thus, for each model and task, we look at the properties of each of the 100 generated texts (20 transmissions chains * 5 seeds) at each generation, which gives us a sample of 100 property values for each value. Using a Kolmogorov–Smirnov test ks_test, we then test whether the sample obtain at each generation comes from the same distribution as the sample obtained after the first generation. If we can confidently reject the hypothesis that the sample of property values at the end of the transmission chain comes from the same distribution as the sample obtained after the first generation, this would confirm that looking at outputs after a single-turn transmission is not enough for predicting output properties in a multi-turn setting.
@@ 3.4 Attractor strength and position
Human cultural evolution shows that cultural traits sometimes evolve towards attractor states, i.e., content that invites convergence even with different starting points kalish_iterated_2007; miton_universal_2015; miton_motor_2020; buskell_what_2017. Therefore, we were interested in whether transmission chains with LLMs would show similar attractor dynamics, and whether these depend on the model and task used in the chain. The concept of cultural attractor is not consistently formalized in the human cultural evolution literature buskell_what_2017. Here, we defined attractors as the theoretical equilibrium point to which the iterated generation process (defined in Eq. 1) may eventually converge. We mathematically define attractors in terms of two properties of interest: its position (i.e., the location in output generation space the process converges toward) and its strength (i.e., the intensity to which generated outputs are pulled toward it). The strength takes values in , which allows for a continuous notion of an attractor: rather than being a binary concept that either exists or does not, attractors here lie on a spectrum, covering systems without attraction effects (strength=0) to ideal attractors (strength=1). To compute position and strength, we use the simulated data to fit a linear regression predicting the value of a property at the end of the chain (i.e. after as a function of its value in the initial text (Figure S1).
@@
where I is the estimated intercept and s the estimated slope. This enables us to estimate the final output of a new chain starting from the final output of the previous chain as:
@@
where . If , then the sequence converges, its limit is and its convergence rate is . We can therefore use the estimated relationship Eq. 2 to determine if an attractor exists () and, if so, estimate its position and strength .
To validate that these theoretical fixed points correctly capture attraction dynamics, we estimated their positions using only data from the first 10 generations of each chain, and compared the predictions with the actual output after 50 generations. Visual inspection of the results (Appendix B) confirmed that our method is suited for estimating the strength and position of attractors.